OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8 @mathslover

OpenStudy (anonymous):

\[Given~\log _{p}4=m~and~\log _{2}Q=n,\]express in terms of m and/or n.

OpenStudy (anonymous):

\[\log _{2}4P^2\]

mathslover (mathslover):

\(\log_2 (2P)^2 \) \(2 \log_2 (2P) \) \(2 [ \log_2 2 + \log_2 P ]\)

mathslover (mathslover):

\(\log_p 4 = m\) \(p^m = 4\) Take log both sides \(\log_2 (p^m) = \log_2 (4) \\ m \log_2 p = 2 \\ \log_2 p = 2/m \)

mathslover (mathslover):

\(\color{blue}{\text{Originally Posted by}}\) @mathslover \(2 [ \log_2 2 + \log_2 P ]\) \(\color{blue}{\text{End of Quote}}\) Plug in the value of \(\log_2 p\) in the above expression. You get : \(2 [ 1 + 2/m ] \)

OpenStudy (anonymous):

\[2+\frac{ 4 }{ m }\]

OpenStudy (anonymous):

Thnx @mathslover :)

mathslover (mathslover):

You're welcome MARC!

OpenStudy (anonymous):

^_^